# Varying Newton constant,entropy and the black hole evaporation law

### v1

@@ Abstract
In Einstein equations we represent the energy-momentum tensor as the one ($T^{\mu\nu}$ ) of an ideal fluid plus the cosmological term. We consider time-dependent Newton “constant” $G$, the cosmological term $\Lambda$ and non-conserved $T^{\mu\nu}$. The Bianchi identity imposes a relation between the energy-momentum (non)conservation and the variation of $G$ and $\Lambda$. If the energy-momentum $T^{\mu\nu}$ is conserved then both the Newton “constant” $G$ and the cosmological term $\Lambda$ either do not change or both must depend on time. If the energy-momentum $T^{\mu\nu}$ is not conserved then the Bianchi identity implies a relation between the energy-momentum and a variation in time of either $G$ or $\Lambda$ (or both). We apply thermodynamics in order to express the non-conservation of the energy-momentum of an ideal fluid by entropy and relate the time variations of $G$ and $\Lambda$ to a change of entropy. Using the relation between a varying Newton constant G and the black hole entropy we derive a modified formula for the Schwarzschild black hole evaporation (a slower evaporation). Its life time is $\frac{9}{5}$ times larger than the one for a constant $G$. The average black hole’s density remains constant when the black hole’s radius shrinks to zero .
@@ 1 Introduction
where $G^{\mu\nu}=R^{\mu\nu}-\frac{1}{2}g^{\mu\nu}R$ $\mu,\nu=0,1,2,3$ and $R^{\mu\nu}$ is the Ricci tensor(we use the conventions of [1]). The form of the energy-momentum tensor $\tilde{T}^{\mu\nu}$ comes from a phenomenological description of fields and matter (we call them fluids). Then, the equations of motion for the fluids can be derived [2][3] from the covariant conservation law of the energy-momentum. In fact, in a wider perspective, the covariant conservation law for $\tilde{T}^{\mu\nu}$ can be applied for a description of fluid dynamics of systems (e.g.,quark-gluon plasma [4])in a flat space with a geometry of a submanifold embedded in the Minkowski space-time .
@@
In this paper we admit varying constants and a non-conservation of the energy-momentum $T^{\mu\nu}$ of an ideal fluid. The violation of the energy-momentum conservation can be a consequence of an incomplete theory ( e.g. some particles (fields) are not taken into account). Some approximations (e.g., diffusion approximation for a particle motion) directly imply a dissipative dynamics. The energy-momentum conservation can be violated if quantum phenomena are included in classical theory. We suggest that the quantum thermodynamic energy conservation (the first law of thermodynamics) can be identified with the conservation of $\tilde{T}^{\mu\nu}$ if we admit a time dependence of $G$ or $\Lambda$ (or both) in the classical Einstein equations (2).
In sec.2 we obtain Bianchi identities resulting from a conservation of $\tilde{T}^{\mu\nu}$(1) and non-conservation of $T^{\mu\nu}$ (2). We show that the conservation of $T^{\mu\nu}$ allows a variation in time of both $G$ and $\Lambda$ but $\partial_{t}G\simeq\partial_{t}\Lambda$. Such models have been discussed in [17][18][19][20].These authors show that a variation of $G$ leads to a variation of $\Lambda$ in an agreement with observations of the universe expansion. In sec.3 applying the thermodynamics we express the variation of cosmological constants by the entropy $S$. According to the first law of thermodynamics $\nabla_{\mu}T^{\mu\nu}$ is related to the heat flow. Then, owing to the second law of thermodynamics the heat flow should be expressed by the change of entropy $TdS$ (where $T$ is the temperature). From the Bianchi identities (applied to eq.(2)) and the first law of thermodynamics it follows that a change of entropy is impossible if $G$ and $\Lambda$ are constant. We discuss a change of $G$ in models with an entropy generated by quantum phenomena at the event horizon. We apply the relation between the entropy and a varying Newton constant $G$ to the Schwarzschild black hole evaporation. As a consequence we obtain a modified evaporation formula. We obtain a slower black hole evaporation. Its life time is $\frac{9}{5}$ times larger than the one for a constant $G$ (that the varying $G$ could change the evaporation law has been discussed in [11][12]). The black hole’s volume tends to zero, but the average energy density defined as a ratio of the mass by the volume does not change during the whole evaporation process ( in contradistinction with the case of a fixed Newton constant). The knowledge of an exact evaporation law may be important for recent searches of the primordial black hole explosions [21][22] raised by an observation of a giant emission of neutrino radiation in KM3NeT [23] (for a search of a final burst of radiation from primordial black holes see [24] and references cited there).
@@ 2 The Bianchi identity
$W=\int dx\sqrt{-g}\Big(R-2\Lambda-8\pi GL\Big),$
where $g$ is the determinant of the metric,$L$ is the Lagrangian for “matter”fields or for an ideal fluid [25][26]. The Bianchi identity applied to eq.(1) leads to the equation
@@
We assume the energy-momentum tensor $T^{\mu\nu}$ in the form of an ideal fluid
$T^{\mu\nu}=(\rho+p)u^{\mu}u^{\nu}+g^{\mu\nu}p,$ (9)
where $\rho$ is the energy density and $p$ is the pressure. The fluid velocity $u^{\mu}$ is normalized as $g_{\mu\nu}u^{\mu}u^{\nu}=-1$. For the ideal fluid we have
$u^{\nu}\nabla_{\mu}T^{\mu}_{\nu}=-(\rho+p)\nabla_{\mu}u^{\mu}-u^{\mu}\partial_{\mu}\rho.$ (10)
As an example of the metric (5) let us consider the FLRW metric describing an expanding isotropic homogeneous universe (with the curvature index k=1,0,1k=-1,0,1)
@@
In the metric (11) in the comoving frame $u=(1,{\bf 0})$ eq.(8) is satisfied if $p$ does not depend on spatial variables as $T^{0}_{j}=0$ and $\nabla_{j}T^{j}_{k}=\partial_{k}p=0$. The fluid velocity satisfies $\nabla_{\mu}u^{\mu}=(-g)^{-\frac{1}{2}}\partial_{\mu}((-g)^{\frac{1}{2}}u^{\mu})=3H$ with $H=a^{-1}\partial_{t}a$. Hence, if $\rho$ and $p$ depend only on time then from eqs.(7) and (9)
@@ 3 Thermodynamics of the moving fluid
On the rhs of eq.(18) we have used the second law of thermodynamics (defining the entropy $S$) saying that there exists an integrating factor (the absolute temperature $T$) such that $dS=T^{-1}dQ$ depends only on the state of the system.
We are going to generalize eq.(18) so that it describes thermodynamics of an ideal fluid moving with a velocity $u^{\mu}$ in a metric more general than in eq.(11). Let us introduce the inverse temperature $\beta$ as a four-vector (for some recent discussions of an inverse four-temperature and the four-entropy, see [31][32][33] and references cited there)
@@
This is the basic equation which we intend to apply for further studies of gravitational systems with a cosmological horizon. In the example of the metric (11) we have (in the comoving coordinates)
$\begin{array}[]{l}\nabla_{\mu}\sigma^{\mu}=u^{\mu}\nabla_{\mu}\sigma+\sigma\nabla_{\mu}u^{\mu}\cr=u^{\mu}\nabla_{\mu}\sigma+3H\sigma=\partial_{t}\sigma+3H\sigma.\end{array}$ (25)
Hence, eq.(24) reads
$8\pi\beta\rho\partial_{t}G+\beta\partial_{t}\Lambda=-8\pi G(\partial_{t}\sigma+3H\sigma)$ (26)
or if we apply eq.(12)
$\partial_{t}\rho+3H{\cal G}=T\partial_{t}\sigma.$ (27)
where ${\cal G}=\rho+p-\sigma T$ is the density of the Gibbs potential. If we set $\rho=\frac{E}{V}$ and $\sigma=\frac{S}{V}$ in eq.(27) then we recover eq.(18).
4 Compact systems with an event horizon
Eq.(24) may have a wider application to space-time phenomena. We have in mind the star contraction [35] or universe evolution ( e.g.,for time-dependent metrics relevant for black holes in an expanding universe [36][37][38]) . In this section we apply an integrated form of eq.(24) to a compact object. We integrate eq.(24) with the measure $d{\bf x}\sqrt{-g}=d{\bf x}\sqrt{|g_{00}|}\sqrt{\gamma}$, where $\gamma$ is the determinant of the spatial metric in eq.(5) ( for a general metric a better way would be to do the space-time integral in the time-interval $(t,t+dt)$ as in [39])
$\int d{\bf x}\sqrt{g}\Big(\beta^{\nu}T^{\mu}_{\nu}\partial_{\mu}G-G\nabla_{\mu}\sigma^{\mu}-\frac{1}{8\pi}\beta^{\nu}\partial_{\nu}\Lambda\Big)=0.$ (28)
We assume that $G$ and $\Lambda$ depend only on time. In the comoving frame from eq.(20) $\beta^{0}=\beta|g_{00}|^{-\frac{1}{2}}$. Calculating the integral (28) of the energy-momentum of an ideal fluid and using $\nabla_{\mu}\sigma^{\mu}=(-g)^{-\frac{1}{2}}\partial_{\mu}((-g)^{\frac{1}{2}}\sigma^{\mu})$ we obtain
$\begin{array}[]{l}\partial_{t}G\int d{\bf x}\sqrt{\gamma}\beta\rho+G\partial_{t}\int d{\bf x}\sqrt{g}\sigma^{0}+G\int d{\bf x}\partial_{j}(\sqrt{\gamma}\sigma^{j})\cr+\frac{1}{8\pi}\int d{\bf x}\sqrt{\gamma}\beta\partial_{t}\Lambda=0.\end{array}$ (29)

### v2

@@ Abstract
In Einstein equations we represent the energy-momentum tensor as the one ($T^{\mu\nu}$ ) of a fluid plus the cosmological term. We consider time-dependent Newton “constant” $G$, the cosmological term $\Lambda$ and non-conserved $T^{\mu\nu}$. The Bianchi identity imposes a relation between the energy-momentum (non)conservation and the variation of $G$ and $\Lambda$. The covariant divergence $\nabla_{\mu}T^{\mu\nu}$ can be related to the first law of thermodynamics. For compact systems of mass $M$ from the Bianchi identity we obtain a power-law relation $G\simeq M^{-\gamma}$ with $\gamma$ depending on pressure or entropy. We discuss radiation and a mass loss described by the Stefan-Boltzmann law. In this formula we insert an expression for the black hole area and its temperature $T$. The Bianchi identity together with a formula for temperature and entropy $S$ determines the index $\gamma$ in the relation between the Newton constant $G$ and the mass $M$. If the entropy $S$ is defined by the equation $dS=T^{-1}dM$ then $\gamma=1$ (the same as for zero pressure). If the formula of Bekenstein-Hawking entropy holds true for time-dependent $G$ then $\gamma=\frac{2}{3}$. We discuss consequences for the evaporation law of some modified expressions for the entropy appearing in effective models of gravity resulting from an interaction with matter fields. In particular, $\gamma=1$ leads to a constant evaporation temperature whereas $\gamma>1$ to a decreasing temperature and luminosity.
@@ 1 Introduction
where $G^{\mu\nu}=R^{\mu\nu}-\frac{1}{2}g^{\mu\nu}R$ $\mu,\nu=0,1,2,3$ and $R^{\mu\nu}$ is the Ricci tensor(we use the conventions of [1]). The form of the energy-momentum tensor $\tilde{T}^{\mu\nu}$ comes from a phenomenological description of fields and matter (we call them fluids). Then, the equations of motion for the fluids can be derived [2][3] from the covariant conservation law of the energy-momentum.
@@
In this paper we admit varying constants and a non-conservation of the energy-momentum $T^{\mu\nu}$ . The violation of the energy-momentum conservation can be a consequence of an incomplete theory ( e.g. some particles (fields) are not taken into account). Some approximations (e.g., diffusion approximation for a particle motion) directly imply a dissipative dynamics. The energy-momentum conservation can be violated if quantum phenomena are included in classical theory. We suggest that the quantum thermodynamic energy conservation (the first law of thermodynamics) can be identified with the conservation of $\tilde{T}^{\mu\nu}$ if we admit a time dependence of $G$ or $\Lambda$ (or both) in the classical Einstein equations (2).
In sec.2 we obtain Bianchi identities resulting from a conservation of $\tilde{T}^{\mu\nu}$(1) and non-conservation of $T^{\mu\nu}$ (2). We show that the conservation of $T^{\mu\nu}$ allows a variation in time of both $G$ and $\Lambda$ but $\partial_{t}G\simeq\partial_{t}\Lambda$. Such models have been discussed in [19][20][21][22].These authors show that a variation of $G$ leads to a variation of $\Lambda$ in an agreement with observations of the universe expansion. In secs.3 and 4 we show that $\nabla_{\mu}T^{\mu\nu}$ can be related to the heat $Q$ according to the formula $dQ=TdS$ where $T$ is the temperature and $S$ is the entropy. We apply this formula to non-equilibrium systems with time-dependent variables. From the Bianchi identities (applied to eq.(2)) and the first law of thermodynamics it follows that a change of entropy is impossible if $G$ and $\Lambda$ are constant. We obtain an equation which relates the time derivatives of $G$ and the mass $M$ of a compact system. If the pressure in the system is zero then this equation implies that $GM$ is constant. In some systems the Bianchi identity leads to a power-law relation $G\simeq M^{-\gamma}$. In sec.5 we discuss a compact system with an event horizon and its mass decrease resulting from a radiation described by the Stefan-Boltzmann law. In such systems the temperature as well as the area of the horizon depend on $G$ and $M$. When the Bianchi identity determines the relation between $G$ and $M$ then the Stefan-Boltzmann law predicts an explicit dependence on time of both $G$ and $M$. We insert the Bekenstein-Hawking formula for the temperature in this radiation law. In a system with zero pressure as a consequence of a constant $GM$ we obtain constant temperature and constant luminosity during the whole evaporation process (no black hole explosion). We derive the same conclusion from the first law of thermodynamics if the entropy is determined by the equation $dS=T^{-1}dM$. We obtain a different result if we assume that Bekenstein-Hawking formula for the entropy applies without any change to models with a time-dependent $G$. In such a case the Bekenstein-Hawking temperature as well as the black hole luminosity tend to infinity at the end of the evaporation process but the mean black hole density remains constant. We discuss the evaporation formula for various possible $\gamma$ in some other models of gravity. In particular, $\gamma>1$ leads to a decreasing black hole temperature and a decreasing luminosity (that the varying $G$ could change the evaporation law has been discussed in [10][11]).
The compact systems with an event horizon could lead to useful tests (besides the ones discussed in [17][18]) of a varying Newton constant $G$ because the area and the temperature of the horizon depend on $G$. The knowledge of an exact evaporation law may be important for recent searches of the primordial black hole explosions [23][24] raised by an observation of a giant emission of neutrino radiation in KM3NeT [25] (for a search of radiation from primordial black holes see [26] and references cited there). A decreasing black hole temperature would be welcome in models of primordial black holes as a cold dark matter.
@@ 2 The Bianchi identity
$W=\int dx\sqrt{-g}\Big(R-2\Lambda-8\pi GL\Big)+K_{GHY},$
where $K_{GHY}$ is the Gibbons-Hawking-York surface term [27][28], $x=(t,{\bf x})$, $g$ is the determinant of the metric,$L$ is the Lagrangian for “matter”fields or for a fluid [29][30]. The Bianchi identity applied to eq.(1) leads to the equation
@@
We consider the energy-momentum tensor $T^{\mu\nu}$ of a fluid
$T^{\mu\nu}=(\rho+p)u^{\mu}u^{\nu}+g^{\mu\nu}p+\Sigma^{\mu\nu},$ (9)
where $\rho$ is the energy density, $p$ is the pressure and $\Sigma^{\mu\nu}$ describes a dissipation (“viscosity” ) of the fluid [32][33] . The fluid velocity $u^{\mu}$ is normalized as $g_{\mu\nu}u^{\mu}u^{\nu}=-1$. We have
$u^{\nu}\nabla_{\mu}T^{\mu}_{\nu}=-(\rho+p)\nabla_{\mu}u^{\mu}-u^{\mu}\partial_{\mu}\rho+u^{\nu}\nabla_{\mu}\Sigma^{\mu}_{\nu}.$ (10)
As an example of the metric (5) let us consider the FLRW solution describing an expanding isotropic homogeneous universe (with the curvature index k=1,0,1k=-1,0,1)
@@
The metric (11) is a solution of Einstein equations with an ideal homogeneous fluid ($\Sigma=0$). In the metric (11) in the comoving frame $u=(1,{\bf 0})$ eq.(8) is satisfied if $p$ does not depend on spatial variables as $T^{0}_{j}=0$ and $\nabla_{j}T^{j}_{k}=\partial_{k}p=0$. The fluid velocity satisfies $\nabla_{\mu}u^{\mu}=(-g)^{-\frac{1}{2}}\partial_{\mu}((-g)^{\frac{1}{2}}u^{\mu})=3H$ with $H=a^{-1}\partial_{t}a$. Hence, if $\rho$ and $p$ depend only on time then from eqs.(7) and (10)
@@
A generalization of eqs.(11)-(12) to $\Sigma\neq 0$ is interesting because it may describe a pasting of FLWR solution with the Schwarzschild solution describing a black hole in an expanding metric [34][35][36][37].
@@ 3 Thermodynamics of the moving fluid
We multiply the Bianchi identity (4) by the velocity $u$, integrate with $\sqrt{-g}d{\bf x}$ and write it in a form showing a relation to the first law of thermodynamics
$\int d{\bf x}\sqrt{-g}\Big(u^{\mu}\partial_{\mu}(G\rho)+G(\rho+p)\nabla_{\mu}u^{\mu}+u^{\nu}\nabla_{\mu}\Sigma^{\mu}_{\nu})\Big)=-\frac{1}{8\pi}\int d{\bf x}\sqrt{-g}u^{\mu}\partial_{\mu}\Lambda,$ (18)
where eq.(10) has been applied. We express eq.(18) in the comoving frame in the metric (5) ( then $g_{00}u^{0}u^{0}=-1$ and $\sqrt{-g}=\sqrt{-g_{00}}\sqrt{\gamma}$, where $\gamma$ is the determinant of the spatial metric). We use
$\nabla_{\mu}u^{\mu}=(\sqrt{-g})^{-1}\partial_{\mu}(u^{\mu}\sqrt{-g})=(\sqrt{-g})^{-1}\partial_{t}\sqrt{\gamma}.$ (19)
Then, eq.(18) reads
$\begin{array}[]{l}\int d{\bf x}\sqrt{\gamma}(\partial_{t}G\rho+G\partial_{t}\rho)+G\int d{\bf x}(\rho+p)\partial_{t}\sqrt{\gamma}\cr=-\int d{\bf x}\sqrt{\gamma}\nabla_{t}\Sigma^{0}_{0}-\frac{1}{8\pi}\int d{\bf x}\sqrt{\gamma}\partial_{t}\Lambda.\end{array}$ (20)
@@
On the rhs of eq.(21) we have used the second law of thermodynamics (defining the entropy $S$) saying that there exists an integrating factor (the absolute temperature $T$) such that $dS=T^{-1}dQ$ depends only on the state of the system. Eq.(21) can be rewritten in the form
$d{\cal H}-Vdp=TdS,$ (22)
where ${\cal H}=E+pV$ is the enthalpy.
It can be seen that the lhs of eq.(20) coincides with the lhs of eq.(21) if we extend the meaning of the differentials of neighboring configurations to configurations at different times. The term $\Sigma$ related to the heat flow in phenomenological thermodynamics is usually also expressed by entropy. We shall assume that it is already contained in $T\partial_{t}S$ and ignore it. Note that if $p=0$ then eq.(20) could be interpreted as the formula (22) for enthalpy with $p$ replaced by $\frac{1}{8\pi}\Lambda$ in the formulation of the black hole pressure according to to ref.[41].
When we define the mass as
$M=\int d{\bf x}\sqrt{\gamma}\rho$ (23)
then eqs.(20)-(21) give
$\begin{array}[]{l}M\partial_{t}G+G(\partial_{t}M+\int d{\bf x}p\partial_{t}\sqrt{\gamma})=-\frac{1}{8\pi}\int d{\bf x}\sqrt{\gamma}\partial_{t}\Lambda\end{array}$ (24)
or in terms of the entropy
$\begin{array}[]{l}M\partial_{t}G+GT\partial_{t}S=-\frac{1}{8\pi}\int d{\bf x}\sqrt{\gamma}\partial_{t}\Lambda.\end{array}$ (25)
In eqs.(23) and (25) we introduced the volume (for the metric (5))
$dV=d{\bf x}\sqrt{\gamma}$ (26)
with $d\partial_{t}V=d{\bf x}\partial_{t}\sqrt{\gamma}$.
For $p=0$ eq.(24) takes a simple form which will be applied in sec.5
$\begin{array}[]{l}\partial_{t}(MG)=-\frac{1}{8\pi}\int d{\bf x}\sqrt{\gamma}\partial_{t}\Lambda.\end{array}$ (27)
There is a long-standing debate over the notion of the energy, volume and the mass in the context of thermodynamics in general relativity [42] (for a recent discussion see [43]). In the classic paper [44] the mass has the ADM meaning [1] of the mass measured at infinity and evaluated on the basis of its gravitational effects. The mass (23) is the mass of the fluid. For relativistic stars eq.(23) agrees with Weinberg’s definition of the mass ( ref.[45], sec.11, eq.(11.1.23)) if $\rho$ is interpreted as the density of nucleons in the star. The mass (23) appears in the formulas for the neutron star. It leads to the relation $G\simeq M^{-\frac{2}{3}}$ ([45], sec.11)in the formula for the Chandrasekhar limt of the neutron star. According to ref.[18] it gives a bound on the variation of $G$. In geometries pasting the FLWR geometry (5) with the Schwarzschild geometry (as discussed in sec.5) these various definitions of the mass may be related.
The reason that we integrated the covariant equation (4) with $d{\bf x}\sqrt{-g}$ is that the space-time volume $dtd{\bf x}\sqrt{-g}$ in eq.(18) does not depend on the choice of coordinates. We could integrate in eq.(18) with $dtd{\bf x}\sqrt{-g}v$ with an arbitrary function $v$. Then, in the integrated form of eq.(24) ( independent of the choice of coordinates if $v$ is a scalar)
$\begin{array}[]{l}\int dt\Big(\tilde{M}\partial_{t}G+G\partial_{t}(\tilde{M}+G\int d{\bf x}vp\partial_{t}\sqrt{\gamma}\Big)\cr=-\frac{1}{8\pi}\int dt\int d{\bf x}\sqrt{\gamma}v\partial_{t}\Lambda\end{array}$
with $\tilde{M}=\int d{\bf x}\sqrt{\gamma}v\rho$. The only change would involve a change of measure $\sqrt{\gamma}\rightarrow v\sqrt{\gamma}$ in the definition of the mass in eq.(23) and the entropy in terms of an entropy density in eq.(25).
We suggest a generalization of eq.(20) to an explicitly covariant formula for relativistic fluids. Let us introduce the inverse temperature $\beta$ as a four-vector (for some recent discussions of an inverse four-temperature and the four-entropy, see [33][46][47] and references cited there)
@@
This is the basic equation which we intend to apply for further studies of gravitational systems with a cosmological horizon. We have $\nabla_{\mu}\sigma^{\mu}=u^{\mu}\nabla_{\mu}\sigma+\sigma\nabla_{\mu}u^{\mu}$. Hence, if we use (4) and (10) then eq.(31) can be rewritten in the form
$u^{\mu}\partial_{\mu}\rho+(\nabla_{\mu}u^{\mu}){\cal G}=\beta^{-1}u^{\mu}\partial_{\mu}\sigma-\beta^{-1}\beta^{\nu}\nabla_{\mu}\Sigma^{\mu}_{\nu}.$ (34)
where ${\cal G}=\rho+p-\beta^{-1}\sigma$ is the density of the Gibbs potential. As (in the comoving frame) $\nabla_{\mu}u^{\mu}=(\sqrt{\gamma})^{-1}\partial_{t}\sqrt{\gamma}\equiv V^{-1}\partial_{t}V$ then (ignoring $\Sigma$)
$\partial_{t}\rho+V^{-1}\partial_{t}V{\cal G}=T\partial_{t}\sigma.$ (35)
When we insert in eq.(35) $\rho=\frac{E}{V}$ and $\sigma=\frac{S}{V}$ then we recover eq.(21). We can also rewrite eq.(35) in the form
$\beta^{-1}\partial_{t}(V\sigma)=\partial_{t}(V\rho)+p\partial_{t}V$ (36)
leading to the correct formula for static systems $\sigma=\beta(\rho+p)$.
4 Compact systems
In this section we show that eq.(31) leads to the same conclusions as eqs.(24)-(25) and (27) of sec.3. We integrate eq.(33) with the measure $d{\bf x}\sqrt{-g}=d{\bf x}\sqrt{|g_{00}|}\sqrt{\gamma}$
$\int d{\bf x}\sqrt{-g}\Big(\beta^{\nu}T^{\mu}_{\nu}\partial_{\mu}G-G\nabla_{\mu}\sigma^{\mu}-\frac{1}{8\pi}\beta^{\nu}\partial_{\nu}\Lambda\Big)=0.$ (37)
We assume that $G$ and $\Lambda$ depend only on time. In the comoving frame $\beta^{0}=\beta|g_{00}|^{-\frac{1}{2}}$ from eq.(28). Calculating the integral (37) of the energy-momentum and using $\nabla_{\mu}\sigma^{\mu}=(-g)^{-\frac{1}{2}}\partial_{\mu}((-g)^{\frac{1}{2}}\sigma^{\mu})$ we obtain
$\begin{array}[]{l}\partial_{t}G\int d{\bf x}\sqrt{\gamma}\beta T^{0}_{0}+G\partial_{t}\int d{\bf x}\sqrt{-g}\sigma^{0}+G\int d{\bf x}\partial_{j}(\sqrt{-g}\sigma^{j})\cr+\frac{1}{8\pi}\int d{\bf x}\sqrt{\gamma}\beta\partial_{t}\Lambda=0.\end{array}$ (38)